OpenStudy (anonymous):

7+3d-5=-1+2d-12+d solve

OpenStudy (anonymous):

collect like terms and simplify..

OpenStudy (anonymous):

having trouble with it

OpenStudy (anonymous):

solve for d do you konw how

OpenStudy (anonymous):

tried solving but I don't think I did it right

OpenStudy (anonymous):

can you show us??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

2+3d=-13+3d +2 +2 3d=-11+3d d=this is where I get lost

OpenStudy (anonymous):

2 + 3d = -13 + 3d subtract 3d from both sides 2 + 3d - 3d = -13 + 3d - 3d combine like terms 2 = -13 (incorrect) There is no solution to this problem

OpenStudy (anonymous):

Solution for 7+3d-5=-1+2d-12+d equation: Simplifying 7 + 3d + -5 = -1 + 2d + -12 + d Reorder the terms: 7 + -5 + 3d = -1 + 2d + -12 + d Combine like terms: 7 + -5 = 2 2 + 3d = -1 + 2d + -12 + d Reorder the terms: 2 + 3d = -1 + -12 + 2d + d Combine like terms: -1 + -12 = -13 2 + 3d = -13 + 2d + d Combine like terms: 2d + d = 3d 2 + 3d = -13 + 3d Add '-3d' to each side of the equation. 2 + 3d + -3d = -13 + 3d + -3d Combine like terms: 3d + -3d = 0 2 + 0 = -13 + 3d + -3d 2 = -13 + 3d + -3d Combine like terms: 3d + -3d = 0 2 = -13 + 0 2 = -13 This equation is invalid, the left and right sides are not equal, therefore there is no solution

OpenStudy (anonymous):

boom

OpenStudy (anonymous):

nicely done jaylan :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

make sure you close the question when your done